OpenStudy (anonymous):

Use the Intermediate Value Theorem to check whether the equation x^3-3x+6.1=0 has a root in the interval (0,1)

OpenStudy (anonymous):

What do you mean the equation has a root? Don't you mean if the function x^3-3x+6.1 has a root?

OpenStudy (anonymous):

Yeah, function

OpenStudy (anonymous):

f(0) = 0^3 - 3*0 + 6.1 = 6.1 f(1) = 1 - 3 + 6.1 = 3.1 So the IVT doesn't tell you anything